### v3

# Gradient-Aware Weight Quantization for Large Language Models

Abstract
Large language models (LLMs) show impressive performance in solving complex language tasks. However, its large number of parameters presents significant challenges for the deployment. So, compressing LLMs to low bits can enable to deploy on resource-constrained devices. To address this problem, we propose gradient-aware weight quantization (GWQ), the first quantization approach for low-bit weight quantization that leverages gradients to localize outliers, requiring only a minimal amount of calibration data for outlier detection. GWQ retains the top 1% outliers preferentially at FP16 precision, while the remaining non-outlier weights are stored in a low-bit. We widely evaluate GWQ on different task include language modeling, grounding detection, massive multitask language understanding and vision-language question and answering. Results show that models quantified by GWQ performs better than other quantization method. During quantization process, GWQ only need one calibration set to realize effective quant. Also, GWQ achieves 1.2$\times$ inference speedup in comparison to the original model and effectively reduces the inference memory.
Gradient-Aware Weight Quantization for Large Language Models
Yihua Shao1,2, Yan Gu3, Siyu Chen3, Haiyang Liu3, Zijian Ling4, Minxi Yan2, Ziyang Yan5, Chenyu Zhang5, Michele Magno6, Haotong Qin6*, Yan Wang3, Jingcai Guo7, Ling Shao1* Hao Tang1* 1PKU  2CASIA  3THU  4USTB  5UNITN  6ETHz  7PolyU  8UCAS thanks: The first two authors contributed equally. correspondence authors: qinhaotong@gmail.com, haotang@pku.edu.cn
1 Introduction
Large language models (LLMs) (Touvron et al., 2023a; Achiam et al., 2023; Almazrouei et al., 2023; Touvron et al., 2023b) based on Transformer (Vaswani, 2017) have demonstrated their outstanding capabilities in scenarios such as language modeling (Brown, 2020). Its ability to handle complex linguistic tasks is due to its large pre-trained data (Kaplan et al., 2020; Yan et al., 2024, 2025) and an astronomical number of parameters (Chowdhery et al., 2023). However, the huge memory consumption leads to great difficulties in its application and deployment (Li et al., 2024a; Remondino et al., 2023). Therefore, when deploying large language models, there are often huge GPU clusters to support the inference of the models (Xu et al., 2024).
In order to apply LLMs to resource-constrained edge devices (Li et al., 2024b; Huang et al., 2024), model compression (Ma et al., 2023; Gu et al., 2023) has become common means to reduce the computational memory. Among them, post-training quantinazation of LLMs (Dettmers and Zettlemoyer, 2023) is a widely used method. Most current quantization methods target to compress modela to 3 or 4 bits (Dettmers et al., 2023; Frantar et al., 2022a), while ensuring models has less loss of performance. Therefore, reducing the size of the model while ensuring its performance is necessary by design effective quantization algorithms (Lin et al., 2024b). With effective quantization algorithms, researchers can make low-bit models’ performance close to that of 16-bit models.
OBQ (Frantar and Alistarh, 2022) indicates that when the pre-trained model has converged, the model should exhibit zero gradients. However, our experiments reveal that the LLMs still generate gradients in response to different text inputs. Inspired by this phenomenon, we propose a gradient-aware post-training weight-only quantization method called GWQ. GWQ is the first post-training quantization approach to utilize gradients to locate outliers in pre-trained models. As same as other quantization method Lin et al. (2024a); Dettmers et al. (2023), GWQ selects 1% sensitive weights as outliers by searching the gradient for the response to the calibration set, preserving these outliers in FP16 precision while quanting the remaining 99% of the weights to 4 bits or 3 bits.
In summary, the primary contributions of this paper are as follows:
  • 1)
    GWQ discovered that locating sensitive weights by first-order gradient is more rational than Hessian matrix.
  • 2)
    GWQ is the first accurate first-order gradient-aware post-training weight quantization method for pre-trained LLMs, requiring only one of calibration data to locating outliers.
  • 3)
    GWQ achieves state-of-the-art (SOTA) in language modeling, grounding detection, massive multitask language understanding and vision-language question and answering tasks in different LLMs. Meanwhile, the quantified models have achieved 1.2$\times$ acceleration compared to the original model and cost less memory during inference.
2 Related Work
Post-training Quantization. Post-training quantization (PTQ) applies mainly to visual models (Gholami et al., 2022). For example, AdaRound (Nagel et al., 2020), AdaQuant (Hubara et al., 2020), and BitSplit (Wang et al., 2024b) can be used for models with less parameters. When models have a large number of parameters, they cause a significant loss but the reduction in memory is not significant. Most quantization methods are developed by Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022). OBQ is developed by the Optimal Brain Surgeon (OBS) (Hassibi et al., 1993; Singh and Alistarh, 2020; Frantar et al., 2021), which default a model response to input is 0 at the end of training. But the OBQ applies weight pruning to model quantization by quantifying the weights that have the less impact on the network. Some PTQ methods locating outlier by reinforcement learning (Wang et al., 2024b) to retrieve the best outlier on a calibration set with a very large sample size. However, the reinforcement learning is not robust and is difficult to converge compared to other PTQ methods, making the calibration process inefficient.
Large Language Model Quantization. Currently, large language model quantization is mainly divided into quantization-based training (Liu et al., 2023; Bondarenko et al., 2024; Zhu et al., 2023; Shao et al., 2025b, a) and post-training quantization. Quantization during training is very difficult to apply due to the large amount of computational resources and overhead required, as well as the large amount of training data. The remaining methods such as ZeroQuant (Yao et al., 2022) utilize knowledge distillation to train one transformer layer at a time instead of training all transformer layers directly. However, poor quality of certain data can also affect the performance of the model after quantization (Wang et al., 2024a). Post-training quantization requires less data compared to quantization-aware training, and only a few thousand or even a few hundred calibration sets are needed to complete the retrieval of sensitive weights for the model. This greatly reduces the data cost and the computational cost of model quantization. Both GPTQ (Frantar et al., 2022a) and SPQR (Dettmers et al., 2023), quantization methods derived from the OBQ method, have shown superior performance in large language models. There are also methods such as AWQ (Lin et al., 2024b), LLM.int8() (Dettmers et al., 2022) that determine the outlier of the model by searching for the activation of the model, and then quantify it with the corresponding method. However, the method of searching for outliers by activation has poor interpretability.
Figure 1: Compared of GWQ and SPQR. Compared to SPQR, the outliers searched by GWQ are much sparser and easier to quantization.
Table 1: Gradient quantization of non-outliers with Hessian in comparison on C4 dataset (Raffel et al., 2020). We localize outliers by gradient and Hessian and quant the non-outliers with RTN (Nagel et al., 2020) to 4 bits.
FP16 (%) base on Gradients FP16 (%) base on Hessian
[…]
Llama-2-7b-hf 6.9759 20.5896 16.8329 15.4319 24.0418 20.2053 18.5902
Llama-2-13b-hf 6.4734 20.3549 16.6689 15.1590 23.9322 20.1040 18.3405
Llama-3-8b-hf 8.9990 22.4648 18.9843 17.5901 22.4423 22.9650 19.3283
3 The Proposed Method
The outlier search method based on Hessian matrix (Frantar et al., 2022a; Dettmers et al., 2023) is built upon the OBQ approach (Frantar et al., 2022b). The OBQ method retrieves the parameters that have the least impact on the model parameters could be expressed as Eq. (1):
$\begin{split}\Delta E=\sum_{i}g_{i}\Delta w_{i}+\frac{1}{2}\sum_{i}h_{ii}\Delta w_{i}^{2}+\\ \frac{1}{2}\sum_{i\neq j}h_{ij}\Delta w_{i}\Delta w_{j}+O\left(\Delta w^{3}\right),\end{split}$ (1)
where $g_{i}=\frac{\partial E}{\partial w_{i}}$ is the first order gradient. In the OBQ framework, it is suggested that once a pre-trained model has fully converged, its gradients $g_{i}=\frac{\partial E}{\partial w_{i}}$ should ideally approach zero.
However, our experiments demonstrate that LLMs continue to generate gradients in response to various text inputs. Bondarenko et al. (2023) denotes that the occurrence of outliers in large models arises from attention heads attempting to avoid updating hidden states. During this process, the softmax function magnifies the formation of strong outliers. Building on this observation, we hypothesize that when a well-trained LLM computes gradients for text input, it often focuses these gradients on irrelevant outliers as a mechanism to prevent hidden state updates. As shown in Fig. 1, compared to SPQR, GWQ demonstrates a sparser allocation of outliers. Specifically, the outliers identified are distributed either per channel or per row, which contrasts with the more dense and uniform distribution of outliers observed in SPQR. This sparsity in GWQ results in a more reasonable allocation, as it does not require concentrating on the compensation of sensitive weight errors across every layer output. As shown in Tab. 1, with different proportions of outliers, the model using Hessian matrix search and quantized using RTN (Nagel et al., 2020) performs worse than the model using gradient search and quantized using RTN on the C4 dataset (Raffel et al., 2020), so it can be learned that it is more reasonable to use the gradient to do outliers localization.
Figure 2: Sensitive weight Location. GWQ utilizes a single calibration sample to calibrate and backpropagates this sample with the subsequent token from the large model’s output as the target label. By halting the backpropagation process, GWQ is able to capture the model gradient $g$ prior to the update of the weights $W$. Within each layer, the gradients associated with each weight block are sorted in relation to their respective weights, and the top 1% with the largest $\left|g\right|$ are identified as the model’s outliers, which are the weights that are particularly sensitive.
3.1 Sensitive Weight Location
GWQ captures the absolute gradients with respect to LLM, acquired through back-propagation, the detail is shown in Alg. LABEL:gradient_alg. We disabled bias in all selected linear layers since the bias term does not get multiplied by inputs in the same way as weights do. This can cause the weight gradients to be unrepresentative of the actual importance of the weights, leading to a disproportionate influence on the MSE loss calculation. As it shown in Fig. 2, the calibration dataset is represented as $D_{c}$, the weights of LLM as $W$, and the loss function of LLM as $\mathcal{L}=(W;D_{c})$. So, the gradients $g$ of the LLM could be expressed as (2),
$g=\bigtriangledown_{W}\mathcal{L}=\bigtriangledown_{W}\mathcal{L}(W;D_{c}),$ (2)
In order to locate the sensitivities as outliers, we calculate loss $\mathcal{L}(W_{Q})$ after model quantization based on Eq. (3).
$\mathcal{L}(W_{Q})\simeq\mathcal{L}(W)-g^{\top}(W-W_{Q}).$ (3)
We define the quantization process as an optimization problem to search for quantitatively optimal outliers. The process can be expressed as Eq. (4),
$Q\left(W\right)=arg\min_{Q}\left\|W-W_{Q}\right\|_{2}^{2},$ (4)
Therefore the final weight optimization formula can be expressed as Eq (5),
$Q\left(W\right)=arg\min_{Q}\bigtriangledown_{W}\mathcal{L}(W;D_{c})(W-W_{Q}).$ (5)
3.2 Gradient Aware Quantization
To quantize the LLM weight $W$, we first compute the scale $s$ and the zero point $z$, which are calculated as depicted in Eq. (6) and (7), where $\beta$ represents the group size (set to 16 in this case) and $b$ is the bit-width for quantization.
$s=\frac{\max(W)-\min(W)}{2^{b-1}},$ (6)
$z=\frac{-\min{W}}{s}.$ (7)
We select the weight by absolute values of the weight gradients $|W_{grad}|$, the largest 1% of them are filtered as outliers, and the rest are used as quantization weights. The process can be expressed as Eq. (8),
$O=\mathbb{I}\left(|W_{grad}|\geq\text{quantile}(\alpha_{1})\right),$ (8)
where $\alpha_{1}$ is the specific quantile value of the magnitude of weight gradients, which are used to identify outliers matrix $O$.
The quantization and dequantization operations are carried out per channel within each group to assess the quantization error. The formulas for quantization and dequantization are provided below (Eq. (9) and (10)), and any identified outliers are excluded from the error computation Eq. (8):
$q=round(\frac{W}{s})+z,$ (9)
Finally, the quantized weight $Q(W)$ can be denoted as Eq (10):
$Q(W)=s\times(q-z).$ (10)
4 Experiments
Table 2: Results of Llama family on language modeling task. GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
Average Calibration Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
[…]
Original 16 - 5.47 / 61.01 6.97 / 56.77 4.88 / 62.88 6.47 / 57.90 6.23 / 58.34 8.99 / 51.71
+ OPTQ 4 1024 9.76 / 51.53 11.24 / 49.01 9.37/ 51.89 10.65 / 50.02 11.56 / 39.34 15.84 /44.93
+ QuIP 4 1024 8.43 / 53.13 10.01 / 50.82 8.13 / 53.48 9.49 / 51.33 10.89 / 40.84 14.03 / 46.22
+ GPTQ 4 1024 5.83 / 59.88 7.79 / 54.74 5.12 / 61.91 7.72 / 56.18 8.21 / 43.78 11.48 / 47.75
+ AWQ 4 512 5.60 / 60.54 7.70 / 55.16 4.97 / 62.53 7.10 / 56.44 6.64 / 57.07 10.69 / 47.75
+ SPQR-O 4.63 1024 5.53 / 60.80 7.02 / 56.63 4.93 / 62.71 6.50 / 57.78 6.42 / 57.70 9.23 / 51.20
+ SPQR-R 3.98 1024 5.55 / 60.56 7.06 / 55.18 4.95 / 62.55 6.53 / 56.45 6.45 / 57.09 9.26 / 48.97
+ GWQ-O (Ours) 4.63 1 5.48 / 60.87 6.99 / 56.72 4.90 / 62.76 6.48 / 53.84 6.40 / 57.83 9.11 / 51.46
+ GWQ-R (Ours) 3.98 1 5.50 / 60.59 7.01 / 56.20 4.38 / 62.56 6.50 / 56.77 6.41 / 57.69 9.15 / 49.89
Table 3: Results of QWEN-VL families on language molding task and grounding detection task. GWQ outperforms other quantization methods on both language modeling task and grounding detection task.
Average Calibration QWEN-VL QWEN-VL-chat
[…]
Original 16 - 8.21 / 54.18 10.10 / 49.51 89.36 9.85 / 52.06 11.87 / 48.07 88.55
+ OPTQ 4 1024 9.34 / 53.13 12.99 /45.76 85.78 13.78 / 50.02 15.71 / 44.93 84.12
+ QuIP 4 1024 9.48 / 53.87 12.68 / 46.41 86.92 12.33 / 51.65 13.98 / 45.76 85.41
+ GPTQ 4 1024 9.31 / 52.07 11.97 / 47.08 87.13 11.50 / 52.06 13.59 / 46.09 86.67
+ AWQ 4 512 8.69 / 53.58 10.98 / 48.12 84.51 11.02 / 50.77 12.88 / 47.43 87.77
+ SPQR-O 4.63 1024 8.54 / 54.05 10.25 / 49.26 88.47 10.13 / 51.99 12.02 / 47.35 88.26
+ SPQR-R 3.98 1024 8.69 / 53.57 10.98 / 48.14 88.00 11.03 / 50.80 12.88 / 46.84 87.78
+ GWQ-O 4.63 1 8.24 / 54.06 10.10 / 49.50 88.98 9.86 / 52.03 11.88 / 48.07 88.49
+ GWQ-R 3.98 1 8.53 / 53.87 10.01 / 49.12 88.42 10.34 / 51.93 12.62 / 47.18 87.93
Figure 3: Performance of Different Quantized Qwen-VL. GWQ performs well compared to other quantization methods in the zero-shot grounding detection task.
Overview. We first apply GWQ to the Llama2 Touvron et al. (2023b) and Llama3 Dubey et al. (2024) family of models and evaluate its feasibility on a language modeling task. We then extend our approach to multimodal models Bai et al. (2023a, b); Liu et al. (2024), measuring the generalizability of our approach to multimodal scenarios with different base models for the grounding detection task, vision language question and answer (VQA). To ensure the fairness of the experiments, we compressed the average number of bits of all the mixed-precision quantized models to less than 4 bits.
4.1 Experiments Setting
Setups. The language-only models we selected include Llama2-7B and 13B (Touvron et al., 2023b) and Llama3-8b Dubey et al. (2024). As for vision-language model, we chose LLaVA-V1.5 Liu et al. (2024), Qwen-VL Bai et al. (2023a, b), Qwen-VL-chat Bai et al. (2023a, b). These models were quantized using the first sample of the RedPajama dataset Weber et al. (2024). To evaluate Perplexity (PPL) and Accuracy (Acc.), we utilize the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) as validation sets. All experiments were conducted on an NVIDIA A800 (80G) GPU.
[…]
$\text{PPL}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(x_{i}\mid x_{0:i-1})\right),$ (11)
where $x_{i}$ represents the predicted token conditioned on the former tokens ($x_{<i}$) in inference step $i$ and $c_{w}$ denotes the average characters per word in the generated sequence.
The next token prediction accuracy can be expressed as Eq. (18),
$\mathrm{Acc}=\frac{N_{\hat{y_{t}}=\overrightarrow{y_{t}}}}{T},$ (12)
where $\hat{y_{t}}$ is the predicted token by LLMs, $\overrightarrow{y_{t}}$ is the ground truth token, $T$ is the number of all tokens.
The exact calculation of next token prediction accuracy will be explained in A.1.
Baselines. Our baseline is categorized into uniform precision quantization and mixed precision quantization. For uniform precision quantization, we choose OPTQ Frantar et al. (2022b), QUIP Chee et al. (2023), GPTQ Frantar et al. (2022a) and AWQ Lin et al. (2024b) as the baseline. For mixed precision quantization, we choose SPQR Dettmers et al. (2023) as the baseline. For fairness, we test the results of SPQR and GWQ with an average bit less than 4 respectively.
4.2 Main Result
We conduct separate evaluations for both pure large language models and large multimodal models. For the large language models, we assess the accuracy of the quantized model in terms of perplexity, zero-shot tasks, and quantized inference memory overhead and latency. In addition to evaluating performance on linguistic tasks for the large multimodal models, we also test the zero-shot target detection task across multiple datasets.
Table 4: Results of Llama family on massive multitask language understanding task. GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
Accuracy (%)
Model Methods
Average Bit
Calib. Num
Average STEM Humanities
Social Sciences
Other
Llama-2-7B-hf Original 16 - 45.96 37.04 43.38 51.84 52.44
+OPTQ 4 1024 33.97 25.34 32.41 38.89 39.22
+QuIP 4 1024 34.58 25.90 33.12 39.02 40.29
+GPTQ 4 1024 36.26 27.95 34.55 40.95 41.59
+AWQ 4 512 38.64 30.13 36.85 43.35 44.24
+SPQR-O 4.63 1024 41.34 33.63 39.38 46.65 47.25
+SPQR-R 3.98 1024 38.66 30.40 36.35 43.57 44.33
+GWQ-O 4.63 1 43.36 34.59 40.94 48.32 50.34
+GWQ-R 3.98 1 39.90 31.89 37.54 45.09 45.09
Llama-2-13B-hf Original 16 - 55.68 44.27 54.43 63.41 60.76
+OPTQ 4 1024 43.32 30.85 39.43 52.59 50.44
+QuIP 4 1024 53.54 31.59 40.22 53.90 51.44
+GPTQ 4 1024 45.47 33.45 41.49 54.46 52.49
+AWQ 4 512 47.35 35.48 43.46 56.1 54.33
+SPQR-O 4.63 1024 49.85 37.59 46.45 58.44 56.93
+SPQR-R 3.98 1024 47.90 35.94 44.23 56.57 54.87
+GWQ-O 4.63 1 52.35 40.82 49.22 61.07 59.34
+GWQ-R 3.98 1 49.04 36.95 45.95 57.35 55.90
Llama-2-13B-hf Original 16 - 65.01 55.39 56.57 76.87 71.35
+OPTQ 4 1024 51.90 43.59 44.49 61.59 57.93
+QuIP 4 1024 54.12 45.93 46.29 64.95 59.35
+GPTQ 4 1024 56.00 47.84 48.35 66.49 61.35
+AWQ 4 512 57.84 49.24 50.43 68.44 63.24
+SPQR-O 4.63 1024 60.84 51.94 52.55 71.44 67.43
+SPQR-R 3.98 1024 58.11 49.96 50.54 68.49 63.46
+GWQ-O 4.63 1 63.12 54.08 55.94 75.33 70.36
+GWQ-R 3.98 1 60.78 51.48 52.46 71.43 66.54
Language Modeling Task. We first evaluate the perplexity and accuracy of the quantized model on the WikiText2 Merity et al. (2016) and C4 Raffel et al. (2020) datasets. As shown in Tab. 2 and Tab. 3, at the same average bits, the GWQ compressed model has lower PPL as well as higher Acc compared to the SPQR compressed model. Such results show that the quantized model with the same compression rate GWQ performs better in language modeling than the quantized model with SPQR.
When the average bit-width in the model after GWQ quantization is lower than 4, it still outperforms the 4-bit full-precision quantized model. This indicates that GWQ quantization is more robust compared to the full-precision approach.
For the Qwen-VL family Bai et al. (2023a, b) of models whose foundational model is not Llama, the quantized model of GWQ can still maintain strong language modeling ability. Thus, it is shown that GWQ can be generalized to different foundational models.
Grounding Detection. The grounding detection tasks are mainly focused on zero-shot target detection. We choose the RefCOCO (Kazemzadeh et al., 2014; Yu et al., 2016) dataset to evaluate the zero-shot capability of the quantized model compared to other quantization methods. Intersection over Union $IoU$, which is denoted as:
$IoU=\frac{Box_{a}\cap Box_{b}}{Box_{a}\cup Box_{b}},$ (13)
where $Box_{b}$ denotes the human-labeled correct result in the dataset labeling (Ground-Truth Box) and $Box_{a}$ denotes the result predicted by the algorithm (Predicted Box).
In the experiment setting, the threshold of $IoU$ is 0.5, so the target detection accuracy Acc. can be defined as:
$Acc.=\frac{correct\_num}{total\_num},$ (14)
where $correct\_num$ is the number of samples detected correctly and $total\_num$ is the total number of samples.
As shown in Tab. 3, Compared to SPQR, GWQ quantized same-precision Qwen-VL family models show a significant increase in detection accuracy. This indicates that the GWQ method is better suited for grounding detection scenarios compared to SPQR. Compared to the full-precision quantization method, the detection ability of the GWQ quantized model is still better than the models quantized by other methods when the average bits of the GWQ quantized model are less than 4. This suggests that the GWQ quantized model not only performs well on language modeling tasks but also has the potential to generalize to multimodal scenarios. The performance is shown in Fig. 3, for the task of visual grounding detection of long context, the GWQ post-quantization model is more advantageous than SPQR post-quantization.
Table 5: Evolution of vison-language question answering task datasets. GWQ outperforms other quantization methods on both LLaVA and Qwen-VL families. Results on ScienceQA dataset, Question classes: NAT is Natural Science, SOC is Social Science, LAN is Language Science, TXT is Text Context, IMG is Image context, NO = No context.
Accuracy (%)
Model Methods
Average Bit
Calib. Num
Average NAT SOC LAN TXT IMG NO
LlaVA-V1.5 Original 16 - 62.94 60.98 63.10 67.27 62.99 63.24 60.04
+OPTQ 4 1024 49.91 48.01 50.12 55.84 48.96 48.92 47.63
+QuIP 4 1024 51.76 49.98 51.92 57.78 50.69 50.63 49.58
+GPTQ 4 1024 52.69 51.01 52.95 58.64 51.73 51.71 50.15
+AWQ 4 512 54.66 53.97 55.37 61.36 54.70 54.66 52.29
+SPQR-O 4.63 1024 58.08 56.09 58.03 62.84 57.88 57.93 55.68
+SPQR-R 3.98 1024 55.23 53.99 55.36 60.36 54.68 54.67 52.31
+GWQ-O 4.63 1 60.55 58.89 60.68 64.98 60.24 60.56 57.94
+GWQ-R 3.98 1 56.99 55.01 57.10 61.82 56.78 56.84 54.36
Qwen-VL-chat Original 16 - 68.89 73.20 65.12 68.77 71.07 66.84 68.35
+OPTQ 4 1024 55.99 55.14 58.46 51.72 54.98 60.76 54.85
+QuIP 4 1024 57.75 56.98 60.02 53.58 56.82 62.34 56.76
+GPTQ 4 1024 58.65 57.99 60.64 54.25 57.84 63.24 57.96
+AWQ 4 512 61.39 60.93 63.32 57.15 60.55 65.90 60.46
+SPQR-O 4.63 1024 64.03 67.56 59.91 63.24 66.25 64.02 63.21
+SPQR-R 3.98 1024 61.38 60.92 63.33 57.14 60.56 65.91 60.43
+GWQ-O 4.63 1 67.39 70.16 62.20 65.91 68.72 66.14 65.85
+GWQ-R 3.98 1 64.03 67.66 59.98 62.12 68.33 63.96 62.14
Qwen-VL Original 16 - 66.70 63.55 66.48 69.02 67.91 66.25 66.99
+OPTQ 4 1024 54.03 50.89 52.97 57.03 56.69 52.98 53.64
+QuIP 4 1024 55.80 52.76 54.81 58.65 58.23 54.94 55.38
+GPTQ 4 1024 56.64 53.57 55.68 59.49 59.07 55.86 56.14
+AWQ 4 512 59.17 56.02 58.32 61.90 61.01 58.72 59.02
+SPQR-O 4.63 1024 61.81 58.64 61.93 64.14 62.43 61.76 61.98
+SPQR-R 3.98 1024 59.17 56.01 58.34 61.92 60.99 58.71 59.04
+GWQ-O 4.63 1 63.87 60.43 63.59 66.34 64.94 63.46 64.46
+GWQ-R 3.98 1 61.37 57.53 60.84 63.74 63.01 61.44 61.87
Table 6: efficiency comparison between different quantization methods. GWQ is not the best compared with other quantization methods.
Average Calibration Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
bit number Speedup ($\uparrow$) Memory ($\downarrow$) Speedup ($\uparrow$) Memory ($\downarrow$) Speedup ($\uparrow$) Memory ($\downarrow$)
Original 16 - $\times$1.00 12.83G $\times$1.00 23.63G $\times$1.00 14.57G
+OPTQ 4 1024 $\times$1.90 3.21G $\times$1.91 6.9G $\times$1.90 3.84G
+QuIP 4 1024 $\times$1.98 3.11G $\times$1.99 6.71G $\times$1.98 3.61G
+GPTQ 4 1024 $\times$2.00 3.11G $\times$2.01 6.71G $\times$2.00 3.61G
+AWQ 4 512 $\times$2.01 3.10G $\times$2.00 6.72G $\times$2.01 3.61G
+SPQR-O 4.63 1024 $\times$1.28 N/A $\times$1.27 N/A $\times$1.28 N/A
+SPQR-R 3.98 1024 $\times$1.26 N/A $\times$1.25 N/A $\times$1.26 N/A
+GWQ-O 4.63 1 $\times$1.23 4.16G $\times$1.22 7.13G $\times$1.23 4.84G
+GWQ-R 3.98 1 $\times$1.22 4.16G $\times$1.21 7.13G $\times$1.22 4.84G
Massive Multitask Language Understanding. In this section, we will discuss the ability of Llama-2 and Llama-3 families using different quantization methods in massive multitask language understanding with MMLU dataset Hendrycks et al. (2021a, b).
As shown in Tab. 5, compared with SPQR, GWQ quantized Llama-2 and Llama-3 families show a substantial improvement in multimodal comprehension accuracy. This indicates that GWQ is more suitable for application in complex scenarios than SPQR. Compared with the full-precision quantization method, when the average bits of the GWQ quantized model are less than 4, the GWQ quantized model still outperforms the models quantized by the other methods in language understanding ability. The results show that our method maintains proficiency in single-task language modeling and also exhibits strong adaptability in complex multi-task scenarios.
Table 7: Effect of the number of calibration set to the quantization. Different numbers of calibration set affect a little to quantifies models’ perform. All calibration set are from Red Pajama1 dataset.
sample_number=1 sample_number=4 sample_number=16 sample_number=64 sample_number=512 sample_number=1024
[…]
Llama2-7B-hf 5.48 6.99 5.49 6.98 5.47 6.98 5.45 6.96 5.46 6.97 5.46 6.96
Llama2-13B-hf 4.90 6.48 4.91 6.47 4.90 6.47 4.89 6.46 4.89 6.48 4.89 6.48
Llama3-8B-hf 6.41 9.15 6.43 9.17 6.42 9.16 6.41 9.15 6.40 9.14 6.40 9.14
Qwen-VL 8.24 10.10 8.23 10.10 8.25 10.11 8.23 10.09 8.23 10.09 8.23 10.09
Qwen-VL-chat 9.86 11.88 9.86 11.88 9.85 11.86 9.85 11.87 9.85 11.87 9.85 11.86
Table 8: Effect of different calibration set to quantization results. different calibration set effects a little to quantization results, but special calibration set of task affects a lot.
WikiText2 C4 Red Pajama1 Red Pajama2 AWQ’s Calibration
[…]
Llama2-7B-hf 5.46 6.96 5.47 6.92 5.48 6.99 5.48 6.98 5.80 7.03
Llama2-13B-hf 4.85 6.47 4.88 6.85 4.90 6.48 4.91 6.47 5.99 6.51
Llama3-8B-hf 6.38 9.14 6.40 9.10 6.41 9.15 6.41 9.16 6.49 9.56
Qwen-VL 8.22 10.08 8.23 10.05 8.24 10.10 8.23 10.11 8.31 10.17
Qwen-VL-chat 9.84 11.89 9.86 11.86 9.86 11.88 9.87 11.87 9.93 11.94
Vision-Language Question Answering. The vision-language question answering (VQA) tasks aim to assess the model’s ability to generate accurate responses to questions based on visual and textual inputs. We evaluate the performance of GWQ and other quantitative methods applied to LlaVA-V1.5 Liu et al. (2024), Qwen-VL Bai et al. (2023a), and Qwen-VL-chat Bai et al. (2023a) models on the ScienceQA dataset Lu et al. (2022).
As shown in Table 4, similar to the previous task, the performance of the GWQ quantized model achieved SOTA on VQA accuracy, compared to the other methods. This suggests that GWQ is highly effective for vision-language question-answering tasks. Even when the average bits of the GWQ quantized models are below 4, the GWQ quantized models still outperform models quantized by other methods. This indicates that the GWQ quantized models not only maintain strong performance in language modeling but also generalize well to complex multimodal tasks like VQA.
Model Efficiency Comparison. In this section, we will discuss the impact of different quantization methods on model acceleration and compression effects.
For the model throughput (tokens/s), the speedup rate can be obtained by calculating the time required to generate 512 tokens after quantization and comparing it with the original model. For the model inference overhead (G), we use the peak memory when the model generates 512 tokens as the maximum inference memory by recording it. As shown in Tab. 6, the model after quantization using GWQ 4bit achieves about 1.2$\times$ of speedup compared to the original model. However, compared to SPQR, the acceleration effect of GWQ is not significant. We believe that the insignificant effect of GWQ is due to the fact that the outlier values saved by GWQ search are sparse, which is not favorable for inference calculation. Compared with the full-precision quantization methods, GWQ and SPQR are not advantageous due to their hardware adaptability.
4.3 Ablation Study
In this section, we test the number of checksets and the effect of different checksets on our quantization method and justify our choice of parameters.
Effect on Number of Calibration. We select different number of samples from the Red Pajama dataset as calibration samples respectively, by taking the gradient mean and searching for the corresponding outlier of the model.
We tested the performance of the model after quantizing different numbers of samples on the WikiText2 and C4 datasets. As shown in Tab. 7, we find that the model effect fluctuates slightly when the number of calibration samples is small, but it does not affect the final quantitative results very much. At the same time, when the number of calibration samples is increased, we find that the effect of the quantized model is similar to that when there is only one sample. Therefore, from the consideration of data efficiency and method robustness, we believe that it is optimal to choose a sample of 1.
Effect on Different Calibration. We choose different calibration data Lin et al. (2024a); Weber et al. (2024); Merity et al. (2016); Raffel et al. (2020) to calibrate the Llama families model Dubey et al. (2024); Touvron et al. (2023b) and test the PPL of the quantized model on the WikiText Merity et al. (2016) and C4 Raffel et al. (2020) datasets.
As shown in Tab. 8, different types of calibration samples do not have a significant effect on the quantized model. We guess that the impact of the calibration data on the quantization is less due to the fact that when collecting the gradients, we used the gradient mean as the outlier indexing basis. However, for specialized calibration sets, such as using the C4 calibration set for quantization calibration, and when testing on the C4 calibration set, the model will show better results. At the same time, we believe that the quantization effect is related to the complexity of the checksum set. When WikiText is used as the check set, its quantized model effect has a slight decrease compared to the rest of the models. This may be due to the fact that the language structure of the WikiText check set is relatively simple, which cannot realize more accurate calibration.
5 Conclusion
This paper introduces GWQ, a weight-only post-training mixed-precision quantization based on first-order gradients. GWQ preserves the weights corresponding to the top 1% of outliers at FP16 bit-width while storing the remaining non-outlier weights in lower-bit. It found that it is easier to obtain more accurate outliers using the first-order gradient than Hessian matrix. GWQ achieves the lowest quantization loss compared to current SOTA methods with only one single calibration sample. In both text-only tasks and multimodel tasks, LLMs quanted by GWQ perform better than other quantization methods, which indecate that GWQ could apply to different models and tasks. Also, GWQ achieves 1.2$\times$ inference acceleration compared to the original model internship and uses less memory in the inference process.
6 Limitation
GWQ gets gradients via back propagation and ranks them accordingly. However, in models such as OPT, the activation function is ReLU, and the correct gradients cannot be computed because of gradient vanishing. Additionally, backpropagation demands substantial memory resources, and a single A100 80G GPU is insufficient to handle the large computational requirements. Furthermore, since GWQ employs mixed-precision quantization, it is less hardware-friendly compared to methods like AWQ, resulting in higher inference latency after quantization. We will continue to refine and optimize the GWQ methodology.
Appendix A Appendix
A.1 Next Token Prediction Accuracy
The accuracy evaluation of for WikiText2 and C4 datasets is based on a causal language modeling task, where given a ground truth sequence of tokens $\mathbf{y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{T}\}$ and the predicted tokens $\hat{\mathbf{y}}=\{\hat{\mathbf{y}}_{1},\hat{\mathbf{y}}_{2},\ldots,\hat{\mathbf{y}}_{T}\}$, the accuracy is computed by shifted ground truth $\vec{\mathbf{y}}$
$\vec{\mathbf{y}}=\mathbf{y}[:,1:]$ (15)
and the most probable token of the model output logits $\mathbf{L}$, where $\mathbf{L}$ is shifted to remove the last token’s predictions
$\mathbf{L}_{\text{shifted}}=\mathbf{L}[:,:-1,:]$ (16)
and predictions $\hat{y}$ are
$\hat{\mathbf{y}}_{t}=\arg\max_{v}\mathbf{L}_{t,v}$ (17)
where $v$ is the vocabulary size. Finally, accuracy is
$\mathrm{Acc}=\frac{N_{\hat{y_{t}}=\overrightarrow{y_{t}}}}{T},$ (18)